Circular No. 7758

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET 2001 W2 S. Nakano, Sumoto, Japan, reports the discovery of a comet by A.

Asami on CCD images taken with the Bisei Spaceguard Center 0.50-m f/2.0 reflector in the course of the “Bisei Asteroid Tracking Telescope for Rapid Survey” program. Additional observations were reported following posting on the NEO Confirmation Page:

2001 UT α_{2000} δ_{2000} m_{1} Observer
Nov. 21.44514 18^{h}11^{m}23.^{s}21 +25^{o}05′20′′.0  Asami
21.44765 18 11 23.73 +25 05 08.4  _{”}
21.45016 18 11 24.15 +25 04 58.9 13.8  _{”}
21.45519 18 11 25.03 +25 04 37.2  _{”}
21.47330 18 11 28.44 +25 03 18.4  _{”}
21.47432 18 11 28.59 +25 03 14.9  _{”}
21.47586 18 11 28.83 +25 03 08.6  _{”}
21.49015 18 11 31.52 +25 02 07.5  _{”}
21.69278 18 12 08.89 +24 47 46.7 11.0 Tichý
21.69840 18 12 09.89 +24 47 22.2  _{”}
21.69975 18 12 10.14 +24 47 16.4  _{”}
21.79939 18 12 28.35 +24 40 12.8  Gajdoš
21.80255 18 12 28.87 +24 39 59.3  _{”}
21.80297 18 12 28.95 +24 39 57.0 14.3  _{”}

A. Asami, D. J. Asher, and N. Hashimoto (Bisei). Coma diameter ∼ 30′′;

tail 30′′ long in p.a. 40^{o}. M. Tichý (Kleť). 0.57-m f/5.2 reflector + CCD. Coma diameter 24′′; 40′′

tail in p.a. 230^{o}; m_{2} = 14.6. Š. Gajdoš and A. Galád (Modra). 0.6-m f/5.5 reflector + CCD.

SUPERNOVA 2001gb IN IC 582 M. Schwartz, Cottage Grove, OR; and W. D. Li, University of Califor-

nia at Berkeley, on behalf of LOTOSS (cf. IAUC 7514), report the discovery of an apparent supernova on an unfiltered image taken with the 0.5-m Tenagra III automated telescope on Nov. 20.5 UT (mag ∼ 16.8). SN 2001gb is confirmed at mag ∼ 18.2 on an earlier image taken by the Katzman Automatic Imaging Telescope (KAIT) on Nov. 11.5. The new object is located at α = 9^{h}59^{m}00^{s}.96, δ = +17^{o}49′12′′.4 (equinox 2000.0), which is ∼ 10′′.8 east and 11′′.1 north of the nucleus of IC 582. A KAIT image taken on Apr. 24.3 showed nothing at this position (limiting mag ∼ 19.0).

2001 November 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2001 CBAT  Daniel W. E. Green